OpenStudy (anonymous):

The odds against Lounging Larry winning the horse race are 3:8. Determine the probability that Lounging Larry wins the horse race. 3/8 8/3 3/11 8/11 5/11

OpenStudy (anonymous):

8/11 - i think!

OpenStudy (anonymous):

ok im not sure how to figure these out..

OpenStudy (anonymous):

3:8 chance of winning=3/(8+3) =3/11